A merchant purchased a line of handbags for R172,00 each. She wants to offer a 14% discount off the listed selling price and still make a 24% profit. The price that she should mark on the label is


1
Expert's answer
2021-05-26T03:29:07-0400

"\text{Cost of the bags is }C=R172,000\\\ \\
\text{She gained a profit of 24\% Profit} =172,000\times\dfrac{24}{100}=1720\times24=R \ 41,280" 

"Sale \ price = Cost +Profit\\
\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =R172,000+R41,280\\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 
=R213,280\\
\therefore Sale \ price = R 213,280\\ \ 
Discount \ given \ on \ sale \ price \ is \ 14\%\\
So, Let \ x \ be \ the \ tag \ price .\\
Since \ the \ discount \ 14\% \ is \ given \ on \ x,\\
So, the \ sale \ price =\dfrac{(100-14)}{100}\times x\\
\Rightarrow \dfrac{86}{100}\times x=213,280\\
\Rightarrow x=\dfrac{(213280\times 100)}{86}=248,000\\\ \\
\therefore The \ price\ tag \ \boxed{x=R 248,000}"